find the mass in grams of 2,053 particles of F2
1
Expert's answer
2020-12-17T07:35:40-0500

6.02 × 1023 particles = 1mole of F2 = 38g F2

1 particle = 38 / (6.02 × 1023 ) g F2

2053 particles = (38 × 2053) / 6.02 × 1023 g F2

= 1.3 × 10-19 g